In​ 2012, the population of a city was 5.82 million. The exponential growth rate was 2.95​% per year.
​a) Find the exponential growth function.
​b) Estimate the population of the city in 2018.
​c) When will the population of the city be ​million?
​d) Find the doubling time.

Answer by Boreal(15235)

F(t)=5.82(1.0295)^t where t= number of years after 2012
F(6) or 2018=5.82^1.0285^6=6,888,882 or 6.89 million
=
1.0295)^t=2
t ln (1.0295)=ln 2
t=23.84 years doubling time
=
c is not clear. For 10 million, 10/5.82=1.718
without rounding
1.718=1.0295^t
t=18.62 years